OpenStudy (anonymous):

A tent is pitched such that it creates an equilateral triangle. If each of the congruent sides measure 2 meters, what is the height of the tent?

OpenStudy (saifoo.khan):

root 3

OpenStudy (saifoo.khan):

(guessed)

OpenStudy (anonymous):

use pythagoras: (1)^2 + x^2 = 2^2 x^2 = 2^2 - 1^2 x^2 = 3 x = sqrt(3)